# Preferential Latent Space Models for Networks with Textual Edges

### v2

@@ Abstract
Many real-world networks contain rich textual information in the edges, such as email networks where an edge between two nodes is an email exchange. The useful textual information carried in the edges is often discarded in most network analyses, resulting in an incomplete view of the relationships between nodes. In this work, we represent each text document as a generalized multi-layer network, and introduce a new and flexible preferential latent space network model that can capture how node-layer preferences directly modulate edge probabilities. We establish identifiability conditions for the proposed model and tackle model estimation with a computationally efficient projected gradient descent algorithm. We further derive the non-asymptotic error bound of the estimator from each step of the algorithm. The efficacy of our proposed method is demonstrated through simulations and an analysis of the Enron email network.
Keywords: latent space model; multi-layer network; non-convex optimization; sparsity; text analysis.
@@ 1 Introduction
Over the past decades, the study of networks has attracted enormous attention, as they naturally characterize complex systems emerging from many research communities, such as social sciences (Bramoullé et al., 2009; Chen et al., 2022), business (Huang et al., 2020; Gualdani, 2021) and economic research (Pesaran and Yang, 2020; Jin et al., 2024). Many statistical network models have been developed to analyze network data, including the exponential random graph model (Frank and Strauss, 1986; Zhang and Chen, 2015), the stochastic block model (Holland et al., 1983), the random dot product model (Athreya et al., 2018), the network-formation model (Leung, 2015; Bräuning and Koopman, 2020), the latent space model (Hoff et al., 2002; Ma et al., 2020), among others.
@@
The majority of statistical research to date has focused on networks with binary (or weighted) edges that characterize the presence (or strength) of connections between nodes. Meanwhile, networks with textual edges, where an edge between two nodes is a text document, are fast emerging. Examples include email networks, where an edge between two nodes representing email accounts is an email, Twitter networks, where an edge between two nodes representing Twitter accounts is a comment or retweet, and coauthor networks, where an edge between two nodes representing authors is a paper. A common approach to analyze such networks is to discard the textual data and use binary or non-negative integer edges to encode the presence or frequency of exchanges. However, this simplified treatment may lose important information encoded in the text data and offer only an incomplete view of the interactions between nodes in the network. For instance, in email networks, discarding the text data carried in edges can overlook important context information in the email exchanges, and is unable to provide insights on the topics communicated in the emails or the sender/receiver preferences of these topics. To gain a more accurate and comprehensive understanding of the relationships between nodes, a more favorable approach is to incorporate the textual information when modeling the formation of network edges.
To extract textual information, useful representation tools from the natural language processing literature can be used. In particular, these tools aim to convert text documents to numerical vectors, which can then be used in text analysis. Some examples include Bag-of-Words that represents tokenized documents as sparse vectors of word counts (Joachims, 1998), sentiment analysis models that convert textual information into sentiment scores (Shapiro et al., 2022), and embedding methods such as Word2Vec (Mikolov et al., 2013), BERT (Devlin et al., 2018), that represent words as dense vectors. These word embedding vectors can capture, for example, semantic similarities and are pre-trained on very large text corpora using neural network models. The appropriate choice of representation tools often depends on the goal of analysis. For example, Word2Vec is useful for learning relationships between words and BERT is useful for translation and next-word prediction (Ke et al., 2023). Dense embedding methods like Word2Vec or BERT may lack interpretability, as it is unclear what each dimension in the embedding space represents, making the vector representations difficult to interpret.
In this paper, we represent each text document between a pair of nodes as a $K$-variate vector that counts the appearances of $K$ keywords in this document. The $K$ keywords are extracted from all text documents using techniques such as textRank (Mihalcea and Tarau, 2004) or topic modeling (Ke et al., 2023). This text representation approach is a form of Bag-of-Words where the dictionary is refined to be keywords in the corpus, and it greatly facilitates model interpretability. Effectively, we convert a text document between nodes $i,j$ to a multivariate edge of dimension $K$, denoted as ${y}_{ijl}=(y_{ijl}^{(1)},\ldots,y_{ijl}^{(K)})$, where $y_{ijl}^{(k)}$ counts the appearance of keyword $k$ in document $l$ between nodes $i,j$. In this paper, we refer to the network with edges ${y}_{ijl}$’s as a generalized multi-layer network, where each edge dimension is referred to as a layer. Unlike a typical multi-layer network, a unique aspect of our data is that the number of nonzero multivariate edges between nodes can vary. For example, nodes $i,j$ may exhange 2 emails, while nodes $i^{\prime},j^{\prime}$ exchange 20. As such, if each email is treated as a separate observation of edges between nodes, our data cannot be conveniently structured into a typical multi-layer network, where no more than one nonzero edge exists between a pair of nodes.
@@
To model this generalized multi-layer network, we propose a new preferential latent space framework that models edge probabilities as a function of latent node positions and node-layer preferences represented via parameter ${W}\in\mathbb{R}^{n\times K}$, where $n$ is the number of nodes in the network and $W_{ik}$ characterizes the interest level of node $i$ on layer $k$. The weight $(W_{ik},W_{jk})$ varies across node pairs and layers, allowing the model to flexibly account for varying interest levels from nodes on layers and give direct insights on how contents of the textual exchanges modulate the relationships between nodes. As the number of keywords $K$ can be large, we further impose a sparsity assumption on ${W}$ to improve model estimability and interpretability. We tackle model estimation with a computationally efficient projected gradient descent algorithm and theoretically derive the error bound of the estimator from each step of the algorithm. Our newly proposed preferential latent space model is general and can be used to model other multi-layer networks, particularly when there are heterogeneous node-layer preferences. Examples include multi-layer social networks, where users interact on different platforms, such as Facebook, Twitter, Instagram, with varying levels of engagement, and international trade networks, where countries trade on different products with varying levels of demands.
@@
Our work contributes to both methodology and theory. As to methodology, we propose a flexible preferential latent space model for generalized multi-layer networks. The proposed model is able to effectively borrow information across a large number of sparse layers when estimating the latent node positions and also provide direct insights into the heterogeneous node-layer preferences. With respect to theory, we establish an explicit error bound for the projected gradient descent iterations that shows an interesting interplay between computational and statistical errors. Specifically, it demonstrates that as the number of iterations increases, the computational error of the estimates converges geometrically to a neighborhood that is within statistical precision of the unknown true parameter. The theoretical analysis is nontrivial, as it involves alternating gradient descent, orthogonal transformation, identifiability constraints, sparsity, and the non-quadratic form of the loss function.
@@
Some recent works considered modeling networks with edges that contain textual information. For example, Sachan et al. (2012); Bouveyron et al. (2018); Corneli et al. (2019); Boutin et al. (2023) considered Bayesian community-topic models that extended the latent Dirichlet allocation model to incorporate network communities. These works assume nodes in the network form several communities and the focus is to identify the community label of each node. Model estimation in these works is often carried out via Gibbs sampling or variational EM, which may be prohibitive when applied to large networks. In comparison, our goal is to understand the relationships between nodes and we do not impose assumptions on the community structure amongst nodes. There is another closely related line of research on modeling standard multi-layer networks, which are special cases of the generalized multi-layer networks we study, by allowing only one nonzero multivariate edge between two nodes. From our data, standard multi-layer networks can be constructed by merging the text documents between a pair of nodes into one. In this case, the edge between nodes $i,j$ on layer $k$ counts the appearance of keyword $k$ in all of the text documents between nodes $i,j$. For multi-layer networks, Paul and Chen (2020); Lei et al. (2020); Jing et al. (2021); Agudze et al. (2022); Lei and Lin (2023); Lyu et al. (2023) and others considered community detection, and Gollini and Murphy (2016); Salter-Townshend and McCormick (2017); D’Angelo et al. (2019) studied Bayesian latent space models. Recently, Zhang et al. (2020); Arroyo et al. (2021) considered multi-layer network models that assume common latent node positions and layer-specific scaling matrices. However, the layer-specific scaling matrix scales all latent node positions equally and is unable to capture the varying level of interests from nodes on a specific layer. Wang et al. (2023) introduced multi-layer random dot product graph model and developed a novel nonparametric change point detection algorithm. MacDonald et al. (2022) proposed a novel latent space model where the latent node positions are concatenations of common position coordinates and layer-specific position coordinates. This model may not work well when there is a large number of sparse layers, as it is challenging to estimate the layer-specific positions in this case. We compare with both Zhang et al. (2020) and MacDonald et al. (2022) in simulations and real data analysis. In particular, we find that our proposed method enjoys better prediction accuracy in the analysis of a real email network.
@@
The rest of our paper is organized as follows. Section 2 introduces the preferential latent space model for networks with multivariate edges and Section 3 discusses model estimation. Section 4 investigates theoretical properties of the estimator from our proposed algorithm. Section 5 reports the simulation results, and Section 6 conducts an analysis of the Enron email corpus data. The paper is concluded with a short discussion section.
@@ 2 Preferential Latent Space Model
Suppose there are $n$ nodes in the network, and between nodes $i,j$, there are $m_{ij}$ document exchanges denoted as $\{{z}_{ijl}\}_{l\in[m_{ij}]}$. Each ${z}_{ijl}$ is a tokenized document consisting of a list of words. From the corpus $\{{z}_{ijl}\}_{i,j\in[n],\,l\in[m_{ij}]}$, we extract a set of $K$ keywords. These keywords can be defined as words with high co-occurrence frequencies with other words in the corpus, using the textRank algorithm (Mihalcea and Tarau, 2004), or as words that represent latent topics within the corpus, using topic modeling techniques (Ke et al., 2023). Correspondingly, each document ${z}_{ijl}$ can be represented as a $K$-dimensional vector ${y}_{ijl}=(y_{ijl}^{(1)},\ldots,y_{ijl}^{(K)})$, where $y_{ijl}^{(k)}$ characterizes the appearance of keyword $k$ in document ${z}_{ijl}$. In this work, we refer to $y_{ijl}^{(k)}$ as an edge when there is no ambiguity and ${y}_{ijl}$ as a multivariate edge. To ease exposition, we focus on undirected binary-valued edges, that is, $y_{ijl}^{(k)}=y_{jil}^{(k)}\in[0,1]$, although our methods and results generalize directly to directed edges and other types of edge values, such as continuous and non-negative integers, using tools in generalized linear models. In summary, the network data we model can be denoted as ${Y}=\{{Y}_{ij}\}_{i,j\in[n]}$, where ${Y}_{ij}\in[0,1]^{m_{ij}\times K}$ collects the $m_{ij}$ multivariate edges between nodes $(i,j)$, and the $l$th row of ${Y}_{ij}$ is the length-$K$ vector, ${y}_{ijl}$. If there is no exchange between nodes $i$ and $j$, we define ${Y}_{ij}=(0,\ldots,0)\in[0,1]^{K}$.
@@
$\text{pr}(Y\mid{U},{\theta})=\prod_{i\neq j,l,k}\text{pr}(y^{(k)}_{ijl}\mid{u}_{i},{u}_{j},{\theta}),$
where ${\theta}$ collects other model parameters to be estimated. Given ${U}$ and ${\theta}$, we assume that $y^{(k)}_{ijl}$ follows a Bernoulli distribution, with $\Theta^{(k)}_{ij}=\text{log odds}(y^{(k)}_{ijl}=1\mid{U},{\theta})$ and
$\Theta^{(k)}_{ij}=a_{i}+a_{j}+(W_{ik}u_{i}^{\top})(W_{jk}u_{j}),$ (1)
where $a_{i}\in\mathbb{R}$ represents the node-specific baseline effect and $W_{ik}\geq 0$ is a weight parameter that quantifies the interest level of node $i$ in layer $k$. In particular, a larger $W_{ik}$ indicates a greater interest of node $i$ in layer $k$. If either node $i$ or $j$ is uninterested in layer $k$, meaning $W_{ik}=0$ or $W_{jk}=0$, then the log odds of $y^{(k)}_{ijl}=1$ reduces to the baseline level $a_{i}+a_{j}$. The parameters ${u}_{i}$ and ${u}_{j}$ are latent node positions, and the angle between them determines the likelihood of edges between nodes $i$ and $j$. When ${u}_{i}$ and ${u}_{j}$ point in the same direction, that is, ${u}_{i}^{\top}{u}_{j}>0$, the two nodes are more likely to have an edge in any layer $k$. Additionally, if both nodes $i$ and $j$ have strong interests in layer $k$, meaning large $W_{ik}$ and $W_{jk}$, the likelihood of an edge between nodes $i$ and $j$ in layer $k$ is further increased. See Figure 1 for an illustration.
@@
We refer to model (1) as the preferential latent space model (PLSM) and discuss model identifiability conditions at the end of this section. In this model, the probability of an edge between nodes $i,j$ in layer $k$ is determined by $W_{ik}W_{jk}{u}_{i}^{\top}{u}_{j}$. The multiplied weight $W_{ik}W_{jk}$ varies across node pairs and layers, allowing the model to flexibly account for varying interest levels from nodes on layers. From this perspective, the proposed model is more flexible than Zhang et al. (2020); Arroyo et al. (2021). The baseline effect ${a}=\left(a_{1},\ldots,a_{n}\right)^{\top}$ and latent positions in ${U}$ are shared across all layers, enabling model (1) to borrow information across a large number of sparse layers when estimating these parameters. When $K$ is large, we impose element-wise sparsity on ${W}$ to ensure its estimability. This stipulates that each node prefers only a subset of the $K$ layers, or equivalently, each layer is only preferred by only a subset of the nodes. This plausible assumption effectively reduces the number of parameters while enhancing model interpretability.
@@
Condition (1) in the above proposition is a norm constraint, imposed to ensure ${W}$ and ${U}$ are identifiable. This condition confines all latent node positions to the unit sphere. Condition (2) assumes at least one entry in ${W}$ is zero and this is imposed to ensure ${W}$ and ${a}$ are identifiable.

### v3

# A Preferential Latent Space Model for Text Networks

@@ Abstract
Network data enriched with textual information, referred to as text networks, arise in a wide range of applications, including email communications, scientific collaborations, and legal contracts. In such settings, both the structure of interactions (i.e., who connects with whom) and their content (i.e., what is communicated) are useful for understanding network relations. Traditional network analyses often focus only on the structure of the network and discard the rich textual information, resulting in an incomplete or inaccurate view of interactions. In this paper, we introduce a new modeling approach that incorporates texts into the analysis of networks using topic-aware text embedding, representing the text network as a generalized multi-layer network where each layer corresponds to a topic extracted from the data. We develop a new and flexible latent space network model that captures how node-topic preferences directly modulate edge formation, and establish identifiability conditions for the proposed model. We tackle model estimation with a projected gradient descent algorithm, and further discuss its theoretical properties. The efficacy of our proposed method is demonstrated through simulations and an analysis of an email network.
Address for correspondence: Jingfei Zhang, Goizueta Business School, Emory University, Atlanta, GA, USA. emma.zhang@emory.edu; Xiaoyue Niu, Department of Statistics, Pennsylvania State University, University Park, PA, USA. xiaoyue@psu.edu.
Keywords: text networks; latent space model; non-convex optimization; text analysis; topic-aware embedding.
@@ 1 Introduction
Over the past decades, the study of networks has attracted enormous attention, as they provide a natural characterization of complex systems emerging from a wide range of research communities, such as social sciences (Borgatti et al., 2009), business (Elliott et al., 2014) and biomedical research (Bota et al., 2003; Zhang et al., 2020). In response to the rising needs in analyzing network data, many statistical network models have been developed, including the exponential random graph model (Frank and Strauss, 1986), the stochastic block model (Wang and Wong, 1987; Nowicki and Snijders, 2001), the random dot product model (Athreya et al., 2018), the latent space model (Hoff et al., 2002; Ma et al., 2020; MacDonald et al., 2022).
@@
The majority of network research to date has focused on networks with binary or weighted edges that characterize the presence or strength of connections between nodes. Meanwhile, networks with textual edges, where an edge between two nodes is a text document, are increasingly common. Examples include email networks, where an edge between two email accounts is an email exchange, and contract networks, where an edge between two firms is a contract. A common approach to analyze such networks is to discard the textual data and use binary or non-negative integer edges to encode the presence or frequency of exchanges. While this simplifies the analysis, it often overlooks important information embedded in the text, such as the topics discussed, the intention of the interaction, or sender/receiver’s preferences. This loss of context can lead to an incomplete or inaccurate understanding of the relationships within the network.
A more principled approach to modeling text networks is to incorporate the textual information when modeling the formation of network edges. Towards this goal, texts can be transformed into numerical representations using tools from natural language processing (NLP). Common techniques include Bag-of-Words (Joachims, 1998), static embedding methods such as Word2Vec (Mikolov et al., 2013), and contextualized embedding methods such as BERT (Devlin et al., 2018) and GPT (Achiam et al., 2023). Modern embedding methods such as BERT and GPT capture rich semantic information and usually perform better in predictive tasks. However, dimensions in the embedding space lack clear interpretation, making the resulting representations difficult to interpret within statistical modeling frameworks.
To form interpretable embeddings for each text document, we consider a topic modeling approach that uses transformer-based embeddings, dimensionality reduction, and clustering to generate interpretable and semantically meaningful topics from text data (Grootendorst, 2022); see details in Section 6. With the extracted topics, we convert each document $l$ between nodes $i$ and $j$ into a multivariate edge of dimension $K$, denoted as ${y}_{ijl}=(y_{ijl}^{(1)},\ldots,y_{ijl}^{(K)})$, where $y_{ijl}^{(k)}\in\{0,1\}$ indicates the presence of topic $k$ in document $l$. We refer to the resulting network as a generalized multi-layer network, where each dimension of the edge corresponds to a distinct topic. This topic-aware embedding greatly enhances model interpretability.
@@
To model this generalized multi-layer network, we propose a new latent space framework that models edge probabilities as a function of latent node positions and node-topic preferences represented via parameter ${W}\in\mathbb{R}^{n\times K}$, where $n$ is the number of nodes in the network and $W_{ik}$ characterizes the interest level of node $i$ on topic $k$. The weight $(W_{ik},W_{jk})$ varies across node pairs and topics, allowing the model to flexibly account for varying interest levels from nodes on topics and give direct insights on how they modulate the relationships between nodes. As the number of topics $K$ can be large, we further impose a sparsity assumption on ${W}$ to improve model estimability and interpretability. We tackle model estimation with a projected gradient descent algorithm and theoretically derive the error bound of the estimator from each step of the algorithm. A particularly useful output of our model is the direct visualization of node positions in a network, offering insights into the varying roles of nodes across different network layers (see Figures 5-6).
@@
In summary, our work contributes to both methodology and theory. As to methodology, we develop a modeling framework for a new and understudied class of network data. We consider topic-aware embedding for text associated with each edge in the network, and propose a flexible multi-layer latent space model. The proposed model is able to effectively borrow information across a large number of sparse layers when estimating the latent node positions and also provide direct insights into the heterogeneous node-topic preferences. With respect to theory, we establish an explicit error bound for the projected gradient descent iterations that shows an interesting interplay between computational and statistical errors. Specifically, it demonstrates that as the number of iterations increases, the computational error of the estimates converges geometrically to a neighborhood that is within statistical precision of the unknown true parameter. The theoretical analysis is nontrivial, as it involves alternating gradient descent, orthogonal transformation, identifiability constraints, sparsity, and a non-quadratic loss function.
@@
Some recent works considered modeling networks with edges that contain textual information. For example, Sachan et al. (2012); Bouveyron et al. (2018); Corneli et al. (2019); Boutin et al. (2023) considered Bayesian community-topic models that extended the latent Dirichlet allocation model to incorporate network communities. These works assume nodes in the network form several communities and the focus is to identify the community label of each node. Model estimation in these works is often carried out via Gibbs sampling or variational EM, which may be prohibitive when applied to large networks. In comparison, our goal is to understand the relationships between nodes and we do not impose assumptions on the community structure amongst nodes. There is another closely related line of research on modeling standard multi-layer networks, which are special cases of the generalized multi-layer networks we study, by allowing only one nonzero multivariate edge between two nodes. From our data, standard multi-layer networks can be constructed by merging the text documents between a pair of nodes into one. In this case, the edge between nodes $i,j$ on layer $k$ counts the appearance of keyword $k$ in all of the text documents between nodes $i,j$. For multi-layer networks, Paul and Chen (2020); Lei et al. (2020); Jing et al. (2021); Agudze et al. (2022); Lei and Lin (2023); Lyu et al. (2023) and others considered community detection, and Gollini and Murphy (2016); Salter-Townshend and McCormick (2017); D’Angelo et al. (2019) studied Bayesian latent space models. Recently, Zhang et al. (2020); Arroyo et al. (2021) considered multi-layer network models that assume layer-specific scaling, but is unable to capture the varying level of interests from nodes on a specific layer. Wang et al. (2023) introduced multi-layer random dot product graph model and developed a novel nonparametric change point detection algorithm. MacDonald et al. (2022) proposed a novel latent space model where the latent node positions are concatenations of common position coordinates and layer-specific position coordinates. This model may not work well when there is a large number of sparse layers, as it is challenging to estimate the layer-specific positions in this case. We compare with both Zhang et al. (2020) and MacDonald et al. (2022) in simulations and real data analysis. In particular, we find that our proposed method enjoys better prediction accuracy in the analysis of a real email network.
@@
The rest of our paper is organized as follows. Section 2 introduces the preferential latent space model for networks with multivariate edges and Section 3 discusses model estimation. Section 4 investigates theoretical properties of the estimator from our proposed algorithm. Section 5 reports the simulation results, and Section 6 conducts an analysis of the Enron email corpus data. The paper is concluded with a discussion section.
@@ 2 Preferential Latent Space Model
Suppose there are $n$ nodes in the network, and between nodes $i,j$, there are $m_{ij}$ document exchanges denoted as $\{{z}_{ijl}\}_{l\in[m_{ij}]}$. Each ${z}_{ijl}$ is a tokenized document consisting of a list of words. From the corpus $\{{z}_{ijl}\}_{i,j\in[n],\,l\in[m_{ij}]}$, we extract a set of $K$ topics (see details in Section 6). Correspondingly, each document ${z}_{ijl}$ can be represented as a $K$-dimensional vector ${y}_{ijl}=(y_{ijl}^{(1)},\ldots,y_{ijl}^{(K)})$, where $y_{ijl}^{(k)}$ characterizes the presence of topic $k$ in document ${z}_{ijl}$. In this work, we refer to $y_{ijl}^{(k)}$ as an edge when there is no ambiguity and ${y}_{ijl}$ as a multivariate edge. We focus on undirected binary-valued edges with $y_{ijl}^{(k)}=y_{jil}^{(k)}\in[0,1]$, although our methods and results generalize directly to directed edges and other types of edge values, such as continuous and non-negative integers, using tools in generalized linear models. We denote the network data we model as ${Y}=\{{Y}_{ij}\}_{i,j\in[n]}$, where ${Y}_{ij}\in[0,1]^{m_{ij}\times K}$ collects the $m_{ij}$ multivariate edges between nodes $(i,j)$, and the $l$th row of ${Y}_{ij}$ is the length-$K$ vector, ${y}_{ijl}$. If there is no exchange between nodes $i$ and $j$, set ${Y}_{ij}=(0,\ldots,0)\in[0,1]^{K}$.
@@
$\text{pr}(Y\mid{U},{\theta})=\prod_{i,j,l,k}\text{pr}(y^{(k)}_{ijl}\mid{u}_{i},{u}_{j},{\theta}),$
where ${\theta}$ collects other model parameters to be estimated. Given ${U}$ and ${\theta}$, we assume that $y^{(k)}_{ijl}$ follows a Bernoulli distribution, with $\Lambda^{(k)}_{ij}=\text{log odds}(y^{(k)}_{ijl}=1\mid{U},{\theta})$ and
$\Lambda^{(k)}_{ij}=a_{i}+a_{j}+(W_{ik}u_{i}^{\top})(W_{jk}u_{j}),$ (1)
where $a_{i}\in\mathbb{R}$ represents the node-specific baseline effect and $W_{ik}\geq 0$ is a weight parameter that quantifies the interest level of node $i$ in topic $k$. In particular, a larger $W_{ik}$ indicates a greater interest of node $i$ in topic $k$. If either node $i$ or $j$ is uninterested in topic $k$, meaning $W_{ik}=0$ or $W_{jk}=0$, then the log odds of $y^{(k)}_{ijl}=1$ reduces to the baseline level $a_{i}+a_{j}$. The parameters ${u}_{i}$ and ${u}_{j}$ are latent node positions, and the angle between them determines the likelihood of edges between nodes $i$ and $j$. When ${u}_{i}$ and ${u}_{j}$ point in the same direction, that is, ${u}_{i}^{\top}{u}_{j}>0$, the two nodes are more likely to have an edge in any topic $k$. Additionally, if both nodes $i$ and $j$ have strong interests in topic $k$, meaning large $W_{ik}$ and $W_{jk}$, the likelihood of an edge between nodes $i$ and $j$ in topic $k$ is further increased. See Figure 1 for an illustration.
@@
We refer to model (1) as the preferential latent space model (PLSM) and discuss model identifiability conditions at the end of this section. In this model, the probability of an edge between nodes $i,j$ in topic $k$ is determined by $W_{ik}W_{jk}{u}_{i}^{\top}{u}_{j}$. The multiplied weight $W_{ik}W_{jk}$ varies across node pairs and topics, allowing the model to flexibly account for varying interest levels from nodes on topics. From this perspective, the proposed model is more flexible than Zhang et al. (2020); Arroyo et al. (2021), where the weight is topic-specific but assumed to be the same across nodes. The baseline effect ${a}=\left(a_{1},\ldots,a_{n}\right)^{\top}$ and latent positions in ${U}$ are shared across all topics, enabling model (1) to borrow information across a large number of sparse layers when estimating these parameters. When $K$ is large, we impose element-wise sparsity on ${W}$ to ensure its estimability. This stipulates that each node prefers only a subset of the $K$ topics, or equivalently, each topic is only preferred by only a subset of the nodes. This plausible assumption effectively reduces the number of parameters while enhancing model interpretability.
@@
(3) ${U},{U}^{\dagger}\in\mathbb{R}^{n\times d}$ have full column rank.
@@
Condition (1) in the above proposition is a norm constraint, imposed to ensure ${W}$ and ${U}$ are identifiable. This condition confines all latent node positions to the unit sphere. Condition (2) assumes at least one entry in ${W}$ is zero and this is imposed to ensure ${W}$ and ${a}$ are identifiable. Condition (3) helps to ensure ${W}$ is identifiable.